# Category-based Galaxy Image Generation via Diffusion Models

### v1

@@ 3.1 Data Collection
The dataset used in this study was derived from the Galaxy10 DECaLS dataset111https://astronn.readthedocs.io/en/latest/galaxy10.html, which contains 17,736 samples with image size pixels colored galaxy images (g, r and z bands) separated in 10 classes. We refined into the Galaxy6 DECaLS dataset, consisting of six categories: Round Smooth Galaxies, In-between Round Smooth Galaxies, Barred Spiral Galaxies, Unbarred Spiral Galaxies, Edge-on Galaxies without Bulges, and Edge-on Galaxies with Bulges. It comprises 14,497 images, distributed across the categories as 2,633, 2,308, 2,035, 4,412, 1,361, and 1,748 images. The dataset was split into training and testing subsets in an 80:20 ratio (Joseph, 2022), providing ample data for the model to learn robust features during training while preserving a representative portion for evaluating its performance on unseen data.
@@
The Galaxy10 DECaLS dataset was constructed from the Galaxy Zoo DR2 (Lintott et al., 2011) and the DESI Legacy Imaging Surveys (Dey et al., 2019) to enable large-scale galaxy classification using high-resolution images. Initially, Galaxy Zoo utilized approximately 270,000 galaxy images from the SDSS dataset, with around 22,000 classified into 10 main categories through volunteer contributions. This dataset was subsequently expanded by incorporating high-quality DECaLS images (Walmsley et al., 2022), resulting in an updated dataset of approximately 441,000 unique galaxies. Through volunteer votes and rigorous filtering, images were categorized into 10 distinct classes: Disturbed Galaxies (1,081 images), Merging Galaxies (1,853), Round Smooth Galaxies (2,645), In-between Round Smooth Galaxies (2,027), Cigar Shaped Smooth Galaxies (334), Barred Spiral Galaxies (2,043), Unbarred Tight Spiral Galaxies (1,829), Unbarred Loose Spiral Galaxies (2,628), Edge-on Galaxies without Bulge (1,423), and Edge-on Galaxies with Bulge (1,873).
@@
3.2 Training Settings
Due to computational resource limitations, the input images in this study were downsampled directly from to , corresponding to the height, width, and number of channels. Through our early experiments, we manually found that reusing the training parameters from previous studies yields the best performance for our model. We set the batch size to 32 and initialized the learning rate to . The Adam optimizer (Kingma, 2014) was employed with hyperparameters and , which are the default parameters in PyTorch222https://pytorch.org/docs/stable/generated/torch.optim.Adam.html.
To enhance training stability, the exponential moving average (EMA) was applied to smooth the model parameters (Tarvainen & Valpola, 2017). The smoothing is computed as
@@ 3.2 Training Settings
The loss function (Eq. 5) in diffusion models primarily reflects the model’s ability to reconstruct noise. However, the quality of generated images is determined by a combination of perceptually relevant attributes, such as semantic consistency, texture details, and overall structure. Thus, the loss value alone is insufficient to comprehensively measure the perceptual quality of the generated images (Dhariwal & Nichol, 2021). To address this limitation, we used the Fréchet Inception Distance (FID; Heusel et al., 2017) to evaluate the model’s generation quality. FID evaluates the quality and diversity of generated samples by comparing the distribution of generated images to that of real images in the feature space of the Inception-V3 (Szegedy et al., 2016). A lower FID score indicates higher generation quality.
@@ 4 Results and Discussion
4.1 Image Quality
As discussed in Section 3.2, we evaluate image quality using the Fréchet Inception Distance (FID). As shown in Table 1, the original Astrddpm shows the least favorable performance. This can primarily be attributed to its inability to capture category variable information. Due to the imbalanced category distribution in the test set and the lack of categorical guidance in Astrddpm, the model struggles to account for quantity discrepancies. As a result, the generated sample distribution deviates from the true test set distribution. Additionally, the significant feature differences between categories make it difficult for the model to generate the correct category-specific features, further amplifying the performance gap. In contrast, when category embeddings are incorporated into Astrddpm, the FID score improves significantly, dropping from 30.5 to approximately 12. This improvement suggests that the inclusion of category embeddings enables the model to effectively capture category-related information, resulting in generated samples that more closely align with the target distribution.
CatGalDiff shows significant improvements in both the quality of individual categories and overall quality. The improvements for each category are 30.02%, 13.82%, 31.81%, 40.2%, 20.95%, and 23.03%, while the overall FID decreases from 11.7 to 5.3, representing an enhancement of 54.39%. In the ablation study of the Astro-RAB, we performed a t-test on the overall FID, resulting in a p-value of 0.35. This indicates that the improvement in image quality due to the Astro-RAB is not statistically significant. In contrast, improvements to the U-Net architecture significantly enhance image generation quality. We attribute this improvement primarily to the additional skip connections (at least two) introduced at the downsampling and upsampling stages in the U-Net. These connections help the decoder preserve high-resolution information, minimizing detail loss and enabling the model to more effectively capture and reconstruct fine structural details. As a result, noise prediction becomes more precise, leading to a significant enhancement in the overall quality of the generated images. We also observed a positive correlation between the improvement rate and the number of instances per category in the dataset, with a Pearson correlation coefficient of 0.715.
@@ 4.1 Image Quality
4.2 Physical properties
To demonstrate that our model captures the measurable features of galaxies, we reference astrddpm and directly compare the galaxy size and flux distributions (Smith et al., 2022). The flux is measured by summing the values within a fixed aperture of 8 pixels in diameter, while the half-light radius represents the size as a simple metric. We use the Wasserstein-1 distance as an indicator to measure the difference between the generated galaxy size and flux distributions and the test set. The Wasserstein-1 distance is a metric for measuring the difference between two probability distributions, and its core idea is to quantify the minimum amount of work required to transform one distribution into another.
For discrete distributions, suppose and are defined on the sets of points and . The Wasserstein-1 distance can be calculated by minimizing the transportation cost. For each pair of points , the transportation cost is given by the absolute distance :
@@ 4.2 Physical properties
We define the sum of the Wasserstein-1 distances between the fluxes in the g-band, r-band, z-band, and the color index (g-r, r-z) of the generated results and the test set as the galaxy color synthetic distance (Color). This can be expressed as:
@@
Then we define the size metric half-light radius Wasserstein-1 distance (), calculated using Eq. 8, and representing the Wasserstein-1 distance between the generated and test set of the half-light radius.
We finally define the weighted synthetic galaxy distance (WSGD) metric as the average of the galaxy color synthetic distance and the half-light radius Wasserstein-1 distance:
This section evaluated the model’s performance using the entire test dataset. Due to the limited sample size and unbalanced category distribution, direct evaluation within each category risked introducing bias, making it insufficient to reflect the model’s generative capabilities. As shown in Table 2, the Astrddpm model consistently performed the worst among all tested methods. By incorporating class embeddings, we improved the model’s ability to represent the physical properties of galaxies, as discussed in Section 4.1. Our Enhanced U-Net architecture demonstrated substantial improvements across all metrics, including a 32.69% increase in galaxy color synthetic distance, a 42.56% improvement in the Wasserstein-1 Distance of the half light radius, and a 39.34% enhancement in WSGD. Adding the Astro-RAB further amplified these gains, achieving a 45.71% improvement in the color metric, a 42.39% increase in the size metric, and a 44.14% enhancement in WSGD. Importantly, the Astro-RAB also achieved a significant reduction in the standard deviation of results, decreasing the standard deviation of the color metric by 42.86%, the size metric by 25.00%, and the WSGD by 45.45%. This indicates that the Astro-RAB significantly improves the consistency of generated results and enhances model stability, further improving the robustness of the model.
@@
As shown in Table 3, the category “Edge-on without Bulge”, which has the fewest samples, performs the worst in terms of the Wasserstein-1 distance for both the color and size metrics and also exhibits the largest standard deviation. The histogram for this category reveals a slight leftward shift in the generated color distribution, deviations in the peak positions for r-band and g-band colors, and a mismatch in the peak of . The category “Barred Spiral,” which has the second-fewest samples in the dataset, performs noticeably better than “Edge-on without Bulge”. The color histograms for this category nearly match the test set’s tail distributions, and the peak positions for g-band, r-band, and closely align with the actual data. However, it is worth noting that the “Round Smooth” category ranks second in terms of differences in the color metric. Upon investigation, we found that the dataset contains several extreme galaxies with exceptionally bright centers with different colors, situated at the right-hand tail of the color histogram. These rare samples are challenging for the diffusion model to capture accurately, leading to generated galaxies that are hard to replicate the color features of these outliers.
Despite some discrepancies in specific categories, the overall distribution of the GalCatDiff model aligns well with the test set, particularly showing significant improvements in the matching of peak positions and long-tail distributions for color and size metrics. This indicates that the model effectively captures global features over the full dataset. However, the model performs poorly in a few categories, characterized by shifts in the tails or mismatches in peak positions of the distributions. This phenomenon can be attributed to the imbalanced category distribution in the dataset (Qin et al., 2023). During training, the model tends to optimize generative performance by learning the data from the dominant categories, thereby capturing the overall feature distribution effectively at a global level. However, because the samples from minority categories contribute less to the overall loss, the model struggles to learn their specific features, resulting in lower generative quality for these categories. In contrast, categories with a larger number of samples contribute more significantly to the loss computation, making the model more attuned to their features. This imbalance leads to inferior generative performance for categories with fewer samples or more complex distributions, as evidenced by deviations in metrics such as color and size.
@@
4.3 Impact of Astro-RAB on Physical Properties
The Enhanced U-Net architecture and the inclusion of the Astro-RAB both make significant contributions to preserving the physical properties of galaxies in the GalCatDiff framework. We performed -tests to compare the WSGD metric improvements: one for the Enhanced U-Net architecture relative to the baseline, and another for the Enhanced U-Net with the Astro-RAB. The resulting -values were and , respectively, both far below the common significance level (e.g., ). The performance improvements of the Enhanced U-Net architecture are consistent with the reasons discussed in Section 4.1, specifically that the additional skip connections help the up-sample stage retain high-resolution information, thereby enhancing the accuracy of noise prediction. Additionally, the improvement in the galaxy color synthetic distance and the significant reduction in the standard deviation across all metrics due to the Astro-RAB, as highlighted in Section 4.2, are particularly remarkable and can be attributed to the structural design of the Astro-RAB.
The core of the Astro-RAB lies in its Attention Fusion Unit design, which integrates convolution and window attention mechanisms. This design not only enables adaptive attention to different regions but also effectively balances the representation of high-frequency and low-frequency components in the image. By leveraging the U-Net to learn weights and output a weighted sum, the module flexibly captures diverse informational features. This capability allows the model to simultaneously address both local and global features, providing significant support for complex galaxy structures and ensuring physical consistency, particularly in galaxy generation tasks.

### v2

@@ III.1 Data Collection
The dataset used in this study was derived from the Galaxy10 DECaLS dataset111https://astronn.readthedocs.io/en/latest/galaxy10.html, which contains 17,736 samples with image size pixels colored galaxy images (g, r and z bands) separated in 10 classes. The original calibrated fluxes in each band are provided in linear units of nanomaggies in the Legacy Surveys222https://www.legacysurvey.org/dr10/description/(Dey et al., 2019). In Galaxy10, the released images are generated by normalizing the nanomaggy fluxes in each band and mapping the resulting g, r, and z values to the 0–255 integer range used for the RGB renderings. We refined into the Galaxy6 DECaLS dataset, consisting of six categories: Round Smooth Galaxies, In-between Round Smooth Galaxies, Barred Spiral Galaxies, Unbarred Spiral Galaxies, Edge-on Galaxies without Bulges, and Edge-on Galaxies with Bulges. It comprises 14,497 images, distributed across the categories as 2,633, 2,308, 2,035, 4,412, 1,361, and 1,748 images. The dataset was split into training and testing subsets in an 80:20 ratio (Joseph, 2022), providing ample data for the model to learn robust features during training while preserving a representative portion for evaluating its performance on unseen data. The detailed counts of images per category in the training, testing, and total sets are reported in Table 1.
| Dataset | Round Smooth | In-between Round Smooth | Barred Spiral | Unbarred Spiral | Edge-on without Bulge | Edge-on with Bulge | Total |
| Train | 2106 | 1846 | 1628 | 3529 | 1088 | 1398 | 11595 |
| Test | 527 | 462 | 407 | 883 | 273 | 350 | 2902 |
| Total | 2633 | 2308 | 2035 | 4412 | 1361 | 1748 | 14497 |
@@
The Galaxy10 DECaLS dataset was constructed from the Galaxy Zoo DR2 (Willett et al., 2013a) and the DESI Legacy Imaging Surveys (Dey et al., 2019) to enable large-scale galaxy classification using high-resolution images. Initially, Galaxy Zoo utilized approximately 270,000 galaxy images from the SDSS dataset, with around 22,000 classified into 10 main categories through volunteer contributions. This dataset was subsequently expanded by incorporating high-quality DECaLS images (Walmsley et al., 2022), resulting in an updated dataset of approximately 441,000 unique galaxies. Through volunteer votes and rigorous filtering, images were categorized into 10 distinct classes: Disturbed Galaxies (1,081 images), Merging Galaxies (1,853), Round Smooth Galaxies (2,645), In-between Round Smooth Galaxies (2,027), Cigar Shaped Smooth Galaxies (334), Barred Spiral Galaxies (2,043), Unbarred Tight Spiral Galaxies (1,829), Unbarred Loose Spiral Galaxies (2,628), Edge-on Galaxies without Bulge (1,423), and Edge-on Galaxies with Bulge (1,873). Although substantially larger datasets are available (e.g., Willett et al., 2013a), in this work we deliberately adopt the refined Galaxy10 DECaLS sample. This decision is motivated by three considerations: (i) limited computational resources, which make it infeasible for us to train diffusion models on extremely large datasets; (ii) the feasibility of conducting additional manual data cleaning at this scale, ensuring higher data quality; and (iii) the relatively balanced number of images across categories, which facilitates more reliable training and evaluation.
@@
After data cleaning, we obtained the final dataset and report the redshift of each galaxy category and its corresponding angular span on the image (pc/pixel), which are presented separately in Fig.4 and Fig.5. It is worth noting that the redshifts were estimated using a random forest algorithm for photometric redshift prediction, and therefore some values may be subject to uncertainties (e.g., Zhou et al., 2021). Moreover, the training and test sets exhibit broadly consistent distributions in both redshift and angular span, with minor deviations in minority categories (e.g., Edge-on without Bulge). This consistency supports the reliability of the test set as a proxy for assessing the model’s generalization ability.
III.2 Training Settings
Due to computational resource limitations, the input images in this study were downsampled directly from to , corresponding to the height, width, and number of channels. The image resolution was thus degraded from to . Through our early experiments, we manually found that reusing the training parameters from previous studies yields the best performance for our model. We set the batch size to 32 and initialized the learning rate to . The Adam optimizer (Kingma, 2014) was employed with hyperparameters and , which are the default parameters in PyTorch333https://pytorch.org/docs/stable/generated/torch.optim.Adam.html. In our training, we did not apply data augmentation strategies. We tested common augmentations such as random rotations and flips, but observed no noticeable improvements, likely because the dataset already contains galaxies with diverse orientations.
To enhance training stability, the exponential moving average (EMA) was applied to smooth the model parameters (Tarvainen and Valpola, 2017). The smoothing is computed as
@@ III.2 Training Settings
The loss function (Eq. 5) in diffusion models primarily reflects the model’s ability to reconstruct noise. However, the quality of generated images is determined by a combination of perceptually relevant attributes, such as semantic consistency, texture details, and overall structure. Thus, the loss value alone is insufficient to comprehensively measure the perceptual quality of the generated images (Dhariwal and Nichol, 2021). To address this limitation, we used the Fréchet Inception Distance (FID; Heusel et al., 2017) to evaluate the model’s generation quality. FID evaluates the quality and diversity of generated samples by comparing the distribution of generated images to that of real images in the feature space of the Inception-V3 (Szegedy et al., 2016). A lower FID score indicates higher generation quality.
@@ IV Results and Discussion
IV.1 Image Quality
As discussed in Section III.2, we evaluate image quality using the Fréchet Inception Distance (FID). For the Total FID reported in Table 2, for the unconditional case (Astrddpm), we generated 2,902 images and compared them against the entire test set of 2,902 real images; for the conditional case (all other models), we used class embeddings to generate images, matching the number of samples in each category to that of the test set. These class-wise generated images were then combined, yielding a total of 2,902 samples, which were compared against the full test set. In both cases, the FID was computed once on the entire generated set versus the entire test set, rather than as a per-class average. Thus, the comparison between the unconditional and conditional models is on equal footing. In addition to the Total FID of Astrddpm, Table 2 also reports the FID for each category, obtained by using class embeddings to generate the same number of samples as in the test set for that category and then computing FID against its test images. As shown in Table 2, the original Astrddpm shows the least favorable performance. This can primarily be attributed to its inability to capture category variable information. Due to the imbalanced category distribution in the test set and the lack of categorical guidance in Astrddpm, the model struggles to account for quantity discrepancies. As a result, the generated sample distribution deviates from the true test set distribution. Additionally, the significant feature differences between categories make it difficult for the model to generate the correct category-specific features, further amplifying the performance gap. In contrast, when category embeddings are incorporated into Astrddpm, the FID score improves significantly, dropping from 30.5 to approximately 12. This improvement suggests that the inclusion of category embeddings enables the model to effectively capture category-related information, resulting in generated samples that more closely align with the target distribution.
CatGalDiff shows significant improvements in both the quality of individual categories and overall quality. The improvements for each category are 30.02%, 13.82%, 31.81%, 40.2%, 20.95%, and 23.03%, while the overall FID decreases from 11.7 to 5.3, representing an enhancement of 54.39%. In the ablation study of the Astro-RAB, we performed a -test on the overall FID, resulting in a p-value of 0.35. This indicates that the improvement in image quality due to the Astro-RAB is not statistically significant. In contrast, improvements to the U-Net architecture significantly enhance image generation quality. We attribute this improvement primarily to the additional skip connections (at least two) introduced at the downsampling and upsampling stages in the U-Net. These connections help the decoder preserve high-resolution information, minimizing detail loss and enabling the model to more effectively capture and reconstruct fine structural details. As a result, noise prediction becomes more precise, leading to a significant enhancement in the overall quality of the generated images. We also observed a positive correlation between the improvement rate and the number of instances per category in the dataset, with a Pearson correlation coefficient of 0.715.
@@ IV.1 Image Quality
IV.2 Physical properties
To demonstrate that our model captures measurable galaxy features, we directly compare the galaxy size and flux distributions of generated and real test images. Both sets are evaluated in the same normalized intensity space using identical preprocessing. Since pixel values are commonly normalized to the [] range for neural-network training stability (e.g., Isola et al., 2017; Zhu et al., 2017; Karras et al., 2019; Ho et al., 2020), the flux values measured from the generated images should be interpreted as proxies for the true physical fluxes, which we refer to as pseudo-flux. Based on the pseudo-flux, we further define a pseudo-mag as an analog of aperture photometry in the g-, r-, and z-bands. Specifically, we measure the pseudo-flux by summing the values within a fixed aperture of 3 pixels in diameter (corresponding to 3 arcsec), and convert it to a pseudo-mag via
This aperture size is also broadly consistent with the 3-arcsec fibers used in the SDSS spectrographs444https://www.sdss4.org/dr12/algorithms/magnitudes/ to define quantities such as fiberMag, and therefore provides a physically meaningful and observationally motivated scale for measuring central fluxes. All color indices reported in this work are computed consistently from the pseudo-mag.
We use the Wasserstein-1 distance as an indicator to measure the difference between the generated galaxy size and flux distributions and the test set. The Wasserstein-1 distance is a metric for measuring the difference between two probability distributions, and its core idea is to quantify the minimum amount of work required to transform one distribution into another. For discrete distributions, suppose and are defined on the sets of points and . The Wasserstein-1 distance can be calculated by minimizing the transportation cost. For each pair of points , the transportation cost is given by the absolute distance :
@@ IV.2 Physical properties
We define a color metric, the galaxy color Wasserstein distance (Color-WD), as the sum of the Wasserstein-1 distances between the pseudo-mag in the g-, r-, and z-bands, and the color index (g-r, r-z) of the generated results and the test set. This can be expressed as:
@@
Before defining the size metric (the half-light radius Wasserstein distance), we describe how the half-light radius is computed. We estimate directly from each normalized galaxy image using a non-parametric procedure consistent with the implementation in the public codebase released by Smith et al. (2022)555https://github.com/Smith42/synthetic-galaxy-distance/. We first measure the median flux within a small central aperture, which provides a stable reference for the central surface brightness. We then expand the radius one pixel at a time and compute the median flux in each newly added annulus, forming a discrete radial brightness profile. The half-light radius is defined as the smallest radius at which the annular median flux drops to 50% of the central value. Because the radius grows in steps of one pixel, the resulting size estimates are inherently discrete. After downsampling the images to resolution, each pixel corresponds to approximately arcsec (Section III.2). Thus, the half-light radii naturally appear quantized at 1-arcsec intervals.
We then define the size metric, half-light radius Wasserstein distance (), calculated using Eq. 9. It is given by
where represents the Wasserstein-1 distance between the generated and test sets of the half-light radius.
We finally define the weighted synthetic galaxy distance (WSGD) metric as the average of the galaxy color Wasserstein distance (Color-WD) and the half-light radius Wasserstein distance ():
This section evaluates the model’s performance using the entire test dataset. While category-wise evaluation is possible, the total test size (2,902 images) and the uneven distribution across categories (see Table 1) mean that some classes contain far fewer examples than others. As a result, per-category metrics may be noisy or biased, so we primarily report overall results, with category-level values provided in Table 3 for reference. As shown in Table 3, the Astrddpm model consistently performed the worst among all tested methods. By incorporating class embeddings, we improved the model’s ability to represent the physical properties of galaxies, as discussed in Section IV.1. Our Enhanced U-Net architecture demonstrated substantial improvements across all metrics, including a 31.65% increase in weighted synthetic galaxy distance, a 41.83% improvement in the Wasserstein-1 Distance of the half light radius, and a 39.77% enhancement in WSGD. Adding the Astro-RAB further amplified these gains, achieving a 25.93% improvement in the color metric, a 43.14% increase in the size metric, and a 38.83% enhancement in WSGD. Importantly, the Astro-RAB also achieved a significant reduction in the standard deviation of results, decreasing the standard deviation of the color metric by 17.67%, the size metric by 26.32%, and the WSGD by 45.45%. In Fig. 7, we present histograms of color metric (g, r, z, color index) and half-light radius for each model on the entire test set (without class separation). Astrddpm remains the weakest performer: it exhibits substantial discrepancies in peak amplitudes (g-band, z-band), fails to align peak positions with the test set (r-band, r–z band, half-light radius), and shows an overall distributional shift (z-band, r–z band). After incorporating class embeddings, the peak positions in the g-band and r-band are corrected; however, the amplitudes change from severe underestimation to slight overestimation, which may indicate mild overfitting. In contrast, our Enhanced U-Net architecture further narrows the gap between the g-band and r-band distributions and improves the peak shape consistency in the z-band and r–z band. After incorporating Astro-RAB, the generated distributions achieved the best agreement with the test set in terms of color index (g–r, r–z) and half-light radius. Together with the improvements in evaluation metrics, this indicates that Astro-RAB improves the consistency of generated distributions with the test set and strengthens model stability, leading to more robust and reliable generative performance.
@@
As shown in Table 4, the category “Edge-on without Bulge”, which has the fewest samples, performs the worst in terms of the Wasserstein-1 distance for both color and size metrics, and shows the largest standard deviation in size. The histogram for this category reveals an overestimation of the g-band peak and shifts in the peak positions of the r-band and z-band colors. These errors further amplify the deviations in the color indices (g–r and r–z). In addition, a mismatch is observed in the peak of the half-light radius distribution. The category “Barred Spiral,” which has the second-fewest samples in the dataset, performs better than “Edge-on without Bulge.” The g- and r-band peaks are correctly located but differ in height from the test set, and the half-light-radius peak is slightly overestimated. However, it is worth noting that the “Round Smooth” category ranks second in terms of differences in the color metric. Upon investigation, we found that the dataset contains several extreme galaxies with exceptionally bright centers with different colors, situated at the right-hand tail of the color histogram. These rare samples are challenging for the diffusion model to capture accurately, leading to generated galaxies that are hard to replicate the color features of these outliers.
Despite some discrepancies in specific categories, the overall distribution of the GalCatDiff model aligns well with the test set, particularly showing significant improvements in the matching of peak positions and long-tail distributions for color and size metrics. This indicates that the model effectively captures global features over the full dataset. However, the model performs poorly in a few categories, characterized by shifts in the tails or mismatches in both the peak positions and peak intensities of the distributions. This phenomenon can be attributed to the imbalanced category distribution in the dataset (Qin et al., 2023). Although we deliberately selected a dataset with relatively balanced categories to mitigate this issue, some discrepancies across classes (e.g., between major and minority categories) remain unavoidable and contribute to the observed differences. During training, the model tends to optimize generative performance by learning the data from the dominant categories, thereby capturing the overall feature distribution effectively at a global level. However, because the samples from minority categories contribute less to the overall loss, the model struggles to learn their specific features, resulting in lower generative quality for these categories. In contrast, categories with a larger number of samples contribute more significantly to the loss computation, making the model more attuned to their features. This imbalance leads to inferior generative performance for categories with fewer samples or more complex distributions, as evidenced by deviations in metrics such as color and size.
@@
IV.3 Impact of Astro-RAB on Physical Properties
The Enhanced U-Net architecture and the inclusion of the Astro-RAB both make significant contributions to preserving the physical properties of galaxies in the GalCatDiff framework. We performed two-sample -tests on the WSGD values obtained from multiple independent runs, taking into account their means and standard deviations. One test compared the Enhanced U-Net architecture against the baseline, and another compared the Enhanced U-Net with and without the Astro-RAB. The resulting -values were and , respectively, both far below the common significance level (e.g., ). The performance improvements of the Enhanced U-Net architecture are consistent with the reasons discussed in Section IV.1, specifically that the additional skip connections help the up-sample stage retain high-resolution information, thereby enhancing the accuracy of noise prediction. Additionally, the improvement in the weighted synthetic galaxy distance and the significant reduction in the standard deviation across all metrics due to the Astro-RAB, as highlighted in Section IV.2, are particularly remarkable and can be attributed to the structural design of the Astro-RAB.
The core of the Astro-RAB lies in its Attention Fusion Unit design, which integrates convolution and window attention mechanisms. Convolutional layers are well-suited for capturing local structures and fine-grained patterns, whereas attention mechanisms excel at modeling long-range dependencies and global context. By leveraging the U-Net to learn weights and output a weighted sum of the convolutional and window-attention features, the module flexibly captures diverse informational features, thereby addressing both local and global characteristics and ensuring physical consistency in the generation of galaxy images.